bruhgetrekt:

\frac{3}{8}x\left(16x^2-32x+24\right)

xxDeppressionxx:

You messed up

bruhgetrekt:

I formatted the question wrong lmao

xxDeppressionxx:

Ik lol

xxDeppressionxx:

Try

Code:
\(\frac{3}{8}x\left(16x^2-32x+24\right)\)

bruhgetrekt:

See that looks a lot better

bruhgetrekt:

How do you make it look like that

xxDeppressionxx:

Code:
\( \frac{3}{8}x\left(16x^2-32x+24\right) \)

xxDeppressionxx:

You put

Code:
\(
at the beginning.

xxDeppressionxx:

And

Code:
\)
at the end

xxDeppressionxx:

Lol

uniqueislove:

:)